OpenStudy (tiffany_rhodes):

Help with an arc length question! There is a telephone wire hanging between two poles x= -b and x= b. It takes the shape of a catenary with the equation y= c + acosh(x/a)

OpenStudy (tiffany_rhodes):

I know the arc length formula is the integral of (1+ f'(t)^2)^(1/2) and the endpoints of integration are -b to b. I just have no idea on how to compute the derivative of y= c+ acosh(x/a). I've never dealt with this function before.

OpenStudy (tiffany_rhodes):

*find the length of the wire.

OpenStudy (anonymous):

\(\cosh{x}\) is the hyperbolic cosine function, defined by \[\cosh{x}=\frac{e^x+e^{-x}}{2}\] It then follows that if \(f(x)=\cosh{x}\), then \[f'(x)=\frac{e^x-e^{-x}}{2}=\sinh{x},\] where \(\sinh{x}\) is the hyperbolic sine function, defined by \[\sinh{x}=\frac{e^x-e^{-x}}{2}.\] In general, then, the derivative of \(a\cosh\dfrac{x}{a}\) is \(a\dfrac{1}{a}\sinh{\dfrac{x}{a}}=\sinh{\dfrac{x}{a}}.\)

OpenStudy (tiffany_rhodes):

Oh okay, so when you square the derivative would it be sin ^2 (x^2/a^2) ?

OpenStudy (tiffany_rhodes):

*sin h^2 (x^2/a^a)

OpenStudy (primeralph):

You don't square the elements in the sinh. You never do that.

OpenStudy (tiffany_rhodes):

hmm. so how would I go about using the arc length formula. I need to square the derivative?

OpenStudy (primeralph):

the derivative, not the elemental "x". Sinh(x)^2 is not sinh(x^2) as you've done.

OpenStudy (tiffany_rhodes):

okay so Sinh(x/a)^2 ?

OpenStudy (primeralph):

That's a part of it....

OpenStudy (tiffany_rhodes):

yeah, I really don't know where I'm going wrong. I've never had to deal with these functions, so I have absolutely no idea what i'm doing.

OpenStudy (primeralph):

Well, flip back to hyperbolics in your textbook. I don't have the time to talk about them. Just know that they follow trig identities similar to sin, cos, tan.

OpenStudy (tiffany_rhodes):

okay.

OpenStudy (primeralph):

Yeah. @sithandwhetever gave some good advice, but may be misleading if you tilt.

OpenStudy (dumbcow):

so plugging into arc length : \[= \int\limits_{-b} ^{b} \sqrt{1+\sinh^{2}(x/a)} dx\] here are hyperbolic identities http://en.wikipedia.org/wiki/Hyperbolic_identities#Useful_relations \[= \int\limits_{-b} ^{b} \sqrt{\cosh^{2} (x/a)} dx\] \[=\int\limits_{-b}^{b} \cosh(x/a) dx\] \[=a \sinh(x/a)\] plug in limits, note symmetry \[= 2a \sinh(b/a)\]

OpenStudy (primeralph):

@dumbcow thank you for taking the time.......

OpenStudy (tiffany_rhodes):

thank you @dumbcow ! :) and everyone else!

OpenStudy (dumbcow):

yw

OpenStudy (primeralph):

Can I get a hug? At least?

OpenStudy (tiffany_rhodes):

lol yes *hug :D

OpenStudy (dumbcow):

haha medals for everyone

OpenStudy (primeralph):

This one?

OpenStudy (tiffany_rhodes):

yes.

OpenStudy (primeralph):

Still don't get it?

OpenStudy (tiffany_rhodes):

so when he said to just plug in the endpoints: 2asinh(b/a) - 2asinh(-b/a) ?

OpenStudy (tiffany_rhodes):

I get the process, just im unsure on the answer

OpenStudy (primeralph):

No that's the final answer.

OpenStudy (tiffany_rhodes):

hmm my math hw says it's wrong. I have no clue.

OpenStudy (primeralph):

Chill.....

OpenStudy (tiffany_rhodes):

i am chill.

OpenStudy (primeralph):

Oh, I didn't mean that in a disrespectful way; I meant hold on

OpenStudy (primeralph):

I'm quite sure 2asinh(b/a) is the answer.

OpenStudy (primeralph):

What's the answer in the book?

OpenStudy (tiffany_rhodes):

It's okay. I'm not using a textbook, it's an online site where I enter my answers

OpenStudy (tiffany_rhodes):

So I have no way to know whether I'm right or not unless it tells me

OpenStudy (primeralph):

did you try typing it as 2*a*sinh(b/a)

OpenStudy (tiffany_rhodes):

yes. It said it was incorrect.

OpenStudy (primeralph):

Okay try asinh(b/a) - asinh(-b/a)

OpenStudy (tiffany_rhodes):

It said it was incorrect again. I can just talk to my math teacher about it tomorrow? She may have made a mistake, it's happened before haha

OpenStudy (primeralph):

I think that's the case.

OpenStudy (primeralph):

One more?

OpenStudy (tiffany_rhodes):

yes, one more problem ;P

OpenStudy (tiffany_rhodes):

I'll just post it on here

OpenStudy (tiffany_rhodes):

Find the arc length function s(x) for the curve (1/3)x^3 +1/(4x) where x>0 with the starting point (1,7/12)

OpenStudy (primeralph):